OpenStudy (anonymous):

Simpilfy the radical expression. -4√160

OpenStudy (anonymous):

\[-4\sqrt{4\times4\times10} = -4\sqrt{4\times4}\sqrt{10} = -4\times4\sqrt{10} = -16\sqrt{10}\]

OpenStudy (anonymous):

Genius.!

OpenStudy (anonymous):

I'm not THAT smart, just a little smart! But thanks anyway!